# PixelWorld: Towards Perceiving Everything as Pixels

### v1

@@ Abstract
Existing foundation models typically process visual input as pixels and textual input as tokens, a paradigm that contrasts with human perception, where both modalities are processed in a unified manner. With the rise of embodied and agentic AI, where inputs primarily come from camera pixels, the need for a unified perception framework becomes increasingly evident. In this paper, we propose to unify all modalities (text, tables, code, diagrams, images, etc) as pixel inputs, i.e. “Perceive Everything as Pixels” (PEAP). We introduce PixelWorld, a novel evaluation suite that unifies all the mentioned modalities into pixel space to gauge the existing models’ performance. Our findings show that (1) PEAP outperforms baseline with token-based input in multimodal datasets, benefiting from unified input for better disambiguation, (2) significant declines in reasoning and coding capabilities across all models when processing pixel-based input, underscoring the need to enhance foundation models’ perceptual abilities, (3) larger models can maintain strong performance on non-reasoning tasks under PEAP, while smaller models like Phi-3.5-V suffer significant performance degradation, (4) the attention pattern of PEAP is highly aligned with text token input, (5) PEAP can be accelerated significantly by exploiting the spatial sparsity. We conclude that the existing frontier models are competent in pixel perception, however, there is still headroom for improvement in the future work.
Machine Learning, Pixel Perception, ICML 2025
@@ 1 Introduction
In recent years, large language models (LLMs) (OpenAI, 2025; Team, 2024) have achieved impressive results across diverse real-world tasks. With advancements in training techniques and model architectures (Brown et al., 2020; Touvron et al., 2023), LLMs process textual data by tokenizing words into subword units and applying causal attention. However, this token-based method differs from how humans and AI agents (e.g., embodied robots (Tellex et al., 2020)) perceive text—namely, through pixel-based visual inputs such as images or screenshots.
This mismatch poses key challenges. First, heuristic pre-processing can be domain- and modality-sensitive (Dagan et al., 2024). Second, morphologically rich or underrepresented languages are often split into disproportionately long sequences (Liang et al., 2023), leading to inefficiencies. Third, even minor typographical errors can significantly degrade performance (Chai et al., 2024). Moreover, crucial format or layout information is usually lost during tokenization. As a result, there is growing interest in vision-language models (VLMs) that directly process text as visual input, potentially improving robustness and preserving spatial or layout cues essential to real-world tasks. This is particularly prominent in areas like embodied agents (Tellex et al., 2020; Driess et al., 2023) and computer-based agents (Zheng et al., 2024; Koh et al., 2024), textual inputs are often captured as pixel-based representations, such as images or video frames obtained via cameras or screenshots.
Although existing foundation models demonstrate strong generalization capabilities, most evaluations in the multimodal domain have primarily focused on image-based semantic understanding, with limited emphasis on understanding and reasoning capabilities for text-rich contexts. This discrepancy highlights the misalignment between current evaluation methodologies and the broader demands of real-world multimodal applications. To bridge this gap, we introduce PixelWorld, an evaluation suite designed to systematically analyze and compare LLMs’ performance on both textual and multimodal tasks when text is perceived as pixel-based input. In Section 2, we describe our dataset collection process, categorizing tasks into three major types: Text-Only, Structural, and Multimodal. For the first two categories, we develop an image synthesis pipeline to transform textual inputs into pixel representations, while for multimodal datasets, we leverage OCR pipeline to extract the text (see Table 1).
In Section 2, we conduct a comprehensive evaluation of models with varying scales on PixelWorld. Sections 3.1 and 3.2 show that applying PEAP to pure-text inputs can degrade performance, but this effect is strongly correlated with task difficulty. No significant decline is observed in simpler tasks such as semantic understanding, whereas more complex tasks like knowledge reasoning and code generation experience substantial drops. This aligns with Insight 2 in Figure 1. Furthermore, larger PEAP models align better with text inputs, while smaller models struggle with instruction following, warranting further investigation. This corresponds to Insight 3 in Figure 1. In Section 3.3, we demonstrate that PEAP excels in intrinsically multimodal tasks, such as website rendering, slide comprehension, and document understanding. Traditional OCR-based preprocessing often introduces errors and information loss, whereas PEAP mitigates these issues, supporting Insight 1 in Figure 1. However, smaller models still exhibit limitations in long-context tasks when using PEAP.
Building on these observations, we provide a deeper discussion in Section 4. In Section 4.1, we visualize the attention patterns of the final layer in Qwen2VL-7B. Despite differences in tokenization (words vs. visual patches), our attention visualizations reveal that PEAP maintains a high degree of consistency in the attention mechanism, correspond to the Insight 4 in the Figure 1. This insight highlights the feasibility of using a vision encoder as a generic universal tokenizer, which could potentially address some of the known issues associated with existing text tokenizers. Furthermore, in Section 4.2, we conduct a quantitative analysis of the potential time overhead using SuperGLUE as an example. We observe that PEAP can lead to increased latency, with processing times up to 3× longer due to larger token sizes. To mitigate this, we implemented a sparsification algorithm, PEAP-Fast, which removes blank pixel regions, significantly accelerating processing times. This approach also aligns more closely with how humans perceive visual content, as empty patches mostly serve to maintain layout (which is already preserved by position embeddings of non-empty patches) and do not warrant additional attention. Section 4.3 evaluates PEAP’s prompt sensitivity, showing that Chain-of-Thought (CoT) prompting boosts performance more effectively than standard methods.
@@
    PixelWorld: This is a comprehensive evaluation suite designed to evaluate pixel-based models across text, structural, and multimodal tasks, enabling direct comparisons between PEAP and token-based approaches.
@@
    Task Performance Analysis: PEAP can enhance structural and multimodal understanding (e.g., website, document) but degrade performance on complex text-centric tasks (e.g., code generation, reasoning). Notably, larger models exhibit better transferability between pixel- and token-based performance, while smaller models struggle with instruction following under PEAP.
@@
    Efficiency and Attention Analysis: We propose PEAP-Fast to optimize the inference speed by removing blank pixel regions, reducing computation overhead without accuracy loss. Additionally, we show that PEAP and token-based models exhibit similar attention patterns, suggesting the possibility of adopting vision encoders as a universal multimodal tokenizer.
@@
GLUE (Wang, 2018) 59,879 Natural language understanding Synthesis test
SuperGLUE (Sarlin et al., 2020) 19,294 Natural language understanding Synthesis test
MMLU-Pro (Wang et al., 2024b) 12,032 Domain knowledge and reasoning Synthesis test
ARC (Clark et al., 2018) 3,548 Science question answering Synthesis test
GSM8K (Cobbe et al., 2021) 1,319 Math problem solving Synthesis test
MBPP (Austin et al., 2021) 757 Programming tasks Synthesis test
@@
TableBench (Wu et al., 2024) 888 Table data understanding and analysis Synthesis test
@@
MathVerse (Zhang et al., 2025) 788 Math and visual reasoning Natural test
MMMU-Pro (Yue et al., 2024) 1,730 Multimodal reasoning Synthesis test
@@ 2 Datasets
Image Data Synthesis
@@ 3 Experiments
Model Selection To validate PixelWorld, we selected a diverse set of vision-language models (VLMs) with varying scales to ensure the robustness and generalizability of our findings. It also allowed us to analyze the behavior of models across different sizes. We evaluated several widely used vision-language models (VLMs), including Qwen2VL-2B (Wang et al., 2024a), Phi-3.5-3.2B (Abdin et al., 2024), Qwen2VL-7B (Wang et al., 2024a), Gemini-Flash (Team, 2024), and GPT-4o (OpenAI, 2025).
@@ 5 Related Work
Multimodal Large Language Models and Benchmarks
Recent progress in multimodal AI has led to the development of models like GPT-4o (OpenAI, 2025), Gemini (Team, 2024), and Claude-3.5 (Anthropic, 2025), which integrate vision-based training to improve instruction-following capabilities. Benchmarks for these models have evolved from task-specific datasets, such as VQA (Agrawal et al., 2016) and DocVQA (Mathew et al., 2021), to more comprehensive evaluations, including MMMU-Pro (Yue et al., 2024), MMBench (Liu et al., 2024), and MegaBench (Chen et al., 2024). However, most current research focuses on the semantic understanding of visual content, with only a few benchmarks—such as MathVerse (Zhang et al., 2025) and MMMU-Pro (Yue et al., 2024)—addressing text recognition and comprehension within images. Our work shifts the focus towards evaluating how well large language models understand language through visual input compared to traditional token-based input.
Screenshot LMs
Recent studies have demonstrated that pretraining on synthetic screenshots can enable vision-language models (VLMs) to achieve performance comparable to that of BERT on language modeling tasks (Lee et al., 2022; Rust et al., 2023; Gao et al., 2024). This approach allows models to better capture text structures without relying on OCR-based methods. Furthermore, our analysis highlights a performance gap between existing VLMs on vision-based tasks and their text-only counterparts, particularly in the absence of relevant pretraining. Interestingly, in certain scenarios, VLMs perform as well as or even better than text-only models, underscoring the potential of this research direction. In the context of document retrieval, recent advancements (Faysse et al., 2024; Ma et al., 2024) have shown that large-scale pretraining on screenshots can outperform traditional OCR-based methods, further reinforcing the advantages of vision-language pretraining.
Language Tokenization
Tokenization methods, such as Byte Pair Encoding (BPE) (Shibata et al., 1999; Sennrich et al., 2016), are widely used in language modeling, but recent studies suggest that they may not always be optimal. For instance, MegaByte (Yu et al., 2023) demonstrated that fixed-length tokenization can improve both computational efficiency and cross-modal capabilities. Similarly, BLT (Pagnoni et al., 2024) proposed entropy-based tokenization, while LCM (team et al., 2024) emphasized the benefits of processing higher-level semantic concepts rather than individual tokens. Inspired by these approaches, we explore whether adaptive image patches can effectively infer textual meaning. At a higher level, we investigate the unification of text and image inputs into a shared representation space, enabling reasoning through abstract semantic concepts rather than traditional token-based methods.
@@ 6 Conclusion
In this work, we introduce PixelWorld, an evaluation suite that unifies text, tables, code, and images into pixel-based inputs to bridge the gap between tokenized text processing and human-like visual perception. Our comprehensive experiments with PEAP show that while pixel-based input improves performance on structurally complex and inherently multimodal tasks (e.g., websites and slides) by reducing OCR errors and preserving contextual layout, it underperforms compared to token-based approaches on challenging text-centric tasks, such as advanced knowledge reasoning and coding. Additionally, attention visualizations reveal a high transferability between pixel patches and textual tokens, indicating the feasibility of future “vision-as-token” approaches. However, pixel-based input incurs higher computational overhead, which we mitigate with PEAP-Fast by removing blank pixel regions, significantly speeding up inference. These findings highlight both the promise and trade-offs of perceiving everything as pixels, offering a unified framework for multimodal understanding while underscoring the need for further research to optimize efficiency and address performance gaps in complex reasoning and coding tasks.

### v2

@@ Abstract
Recent agentic language models increasingly need to interact directly with real-world environments containing intertwined visual and textual information through raw camera pixels, rather than relying on separate image and tokenized text processing, underscoring the necessity of a unified perception paradigm. To close this gap, we explore this idea through Perceive Everything as Pixels (PEAP) and release PixelWorld, a benchmark that renders natural-language, tabular, mathematical and diagrammatic inputs into a single pixel space. Experiments show that PEAP attains competitive accuracy on semantic-understanding tasks, indicating that a vision transformer can capture global textual semantics without explicit tokens. In contrast, reasoning-intensive benchmarks (math and code) exhibit sharp performance drops; however, Chain-of-Thought prompting partially mitigates this gap, hinting that explicit reasoning traces compensate for the missing token structure. We also find that when visual and textual information are closely integrated, representing everything as pixels reduces preprocessing complexity and avoids misalignment issues that often arise in separate pipelines. PixelWorld therefore serves as a practical benchmark for evaluating unified vision–language models and supports broader exploration of PEAP across diverse tasks.
PixelWorld: Towards Perceiving Everything as Pixels
Zhiheng Lyu1,2 Xueguang Ma1 Wenhu Chen1,2 1University of Waterloo  2Vector Institute, Toronto {z63lyu,x93ma,wenhuchen}@uwaterloo.ca
@@ 1 Introduction
In recent years, large vision-language models (L-VLMs) (Wang et al., 2024a; OpenAI, 2025; Team, 2024) have achieved impressive performance across a wide range of real-world tasks. These models typically process visual inputs as pixels and textual inputs as discrete tokens—two distinct modalities that are handled separately. However, such modality-specific processing leads to a fragmented understanding of multimodal inputs and increases the complexity of engineering pipelines. This separation becomes particularly problematic in modern agent-based systems such as Computer Agents (Zheng et al., 2024; Koh et al., 2024) and Embodied Agents (Tellex et al., 2020; Driess et al., 2023), which are increasingly expected to perform complex real-world tasks including navigation in physical environments (Elnoor et al., 2024), booking flights (Chen et al., 2024a), and repairing software bugs on platforms like GitHub (Yang et al., 2024). These tasks involve deeply intertwined visual and textual information, where decoupled tokenization and perception modules can result in high preprocessing overhead (Xie et al., 2024; Koh et al., 2024) and degraded performance due to information loss and layout inconsistencies (Dagan et al., 2024; Chai et al., 2024).
Due to these limitations, we propose a unified perception paradigm: Perceive Everything as Pixels (PEAP). In this paradigm, both text and visual inputs are treated uniformly in the pixel space, allowing a vision-language model (VLM) to jointly model multimodal inputs without separate tokenization or modality-specific encoders. To identify the benefits and challenges of this paradigm, we introduce PixelWorld, a comprehensive benchmark suite designed to assess how well VLMs perform on existing benchmarks under the PEAP setting.
In PixelWorld, we select 10 representative commonly used benchmarks, covering a diverse range of modalities and task scenarios. For each dataset, we construct both traditional token-based and pixel-based (PEAP) input formats using image synthesis and OCR techniques (see Table 1). We then evaluate vision–language models of varying scales, from Qwen2VL-2B to GPT-4o. Cross-modal evaluation in Section 2 yields three overarching insights: Insight 1: in intrinsically multimodal settings such as website rendering, slide comprehension, and document understanding, PEAP eliminates OCR noise and consistently boosts performance; Insight 2: pixelising inputs for reasoning intensive tasks such as math and code incurs marked accuracy drops, yet the gap narrows as model capacity grows, suggesting scale is critical for cross-modal transfer; Insight 3: larger models demonstrate superior instruction-following and long-context reasoning across modalities, while smaller models struggle, emphasizing the need for scale-aware training under the pixel paradigm.
To further understand these findings, we conduct additional analyses from three perspectives: (1) Representation analysis: We visualize the attention patterns of Qwen2VL-7B and find consistent global structures between token- and pixel-based inputs, suggesting that vision encoders can serve as universal tokenizers. (2) Efficiency optimization: We measure inference latency and show that while PEAP increases processing time due to input size, our proposed PEAP-Fast algorithm prunes blank patches and achieves up to 80% speedup without accuracy loss. (3) Prompt sensitivity: We explore input prompting strategies and find that Chain-of-Thought (CoT) boosts performance on PEAP more effectively than standard input.
@@
    PixelWorld: We introduce a unified benchmark that converts text, structural, and multimodal datasets into pixels, providing a direct stress test and diagnostic tool for PEAP. The dataset and evaluation code are publicly released to foster a more holistic yardstick for future L-VLMs research and to ease multimodal data collection.
@@
    Task–scale insights: PEAP consistently improves layout-heavy or intrinsically multimodal tasks (e.g., website and document understanding) but degrades on reasoning- and code-centric benchmarks; the performance gap diminishes as model size increases, highlighting scale as a key factor for transferability.
@@
    Efficiency & interpretability: We propose PEAP-Fast, which removes blank pixel patches to achieve up to a $3\times$3 × latency reduction without harming accuracy. Attention visualizations reveal similar global patterns between pixel- and token-based models, suggesting that vision encoders can act as a universal multimodal tokenizer.
@@
GLUE Wang (2018) 59,879 Natural language understanding Synthesis test
SuperGLUE Sarlin et al. (2020) 19,294 Natural language understanding Synthesis test
MMLU-Pro Wang et al. (2024b) 12,032 Domain knowledge and reasoning Synthesis test
ARC Clark et al. (2018) 3,548 Science question answering Synthesis test
GSM8K Cobbe et al. (2021) 1,319 Math problem solving Synthesis test
MBPP Austin et al. (2021) 757 Programming tasks Synthesis test
@@
TableBench Wu et al. (2024) 888 Table data understanding and analysis Synthesis test
@@
MathVerse Zhang et al. (2025) 788 Math and visual reasoning Natural test
MMMU-Pro Yue et al. (2024) 1,730 Multimodal reasoning Synthesis test
@@ 2 Datasets
2.0.0.0.1 Image Data Synthesis
@@ 3 Experiments
Model Selection To validate PixelWorld, we selected a diverse set of vision-language models (VLMs) with varying scales to ensure the robustness and generalizability of our findings. It also allowed us to analyze the behavior of models across different sizes. We evaluated several widely used vision-language models (VLMs), including Qwen2VL-2B Wang et al. (2024a), Phi-3.5-3.2B Abdin et al. (2024), Qwen2VL-7B Wang et al. (2024a), Gemini-Flash Team (2024), and GPT-4o OpenAI (2025).
@@ 5 Related Work
5.0.0.0.1 Multimodal Large Language Models and Benchmarks
Recent progress in multimodal AI has led to the development of models like GPT-4o OpenAI (2025), Gemini Team (2024), and Claude-3.5 Anthropic (2025), which integrate vision-based training to improve instruction-following capabilities. Benchmarks for these models have evolved from task-specific datasets, such as VQA Agrawal et al. (2016) and DocVQA Mathew et al. (2021), to more comprehensive evaluations, including MMMU-Pro Yue et al. (2024), MMBench Liu et al. (2024), and MegaBench Chen et al. (2024b). However, most current research focuses on the semantic understanding of visual content, with only a few benchmarks—such as MathVerse Zhang et al. (2025) and MMMU-Pro Yue et al. (2024)—addressing text recognition and comprehension within images. Our work shifts the focus towards evaluating how well large language models understand language through visual input compared to traditional token-based input.
5.0.0.0.2 Screenshot LMs
Recent studies have demonstrated that pretraining on synthetic screenshots can enable vision-language models (VLMs) to achieve performance comparable to that of BERT on language modeling tasks Lee et al. (2022); Rust et al. (2023); Gao et al. (2024). This approach allows models to better capture text structures without relying on OCR-based methods. Furthermore, our analysis highlights a performance gap between existing VLMs on vision-based tasks and their text-only counterparts, particularly in the absence of relevant pretraining. Interestingly, in certain scenarios, VLMs perform as well as or even better than text-only models, underscoring the potential of this research direction. In the context of document retrieval, recent advancements Faysse et al. (2024); Ma et al. (2024) have shown that large-scale pretraining on screenshots can outperform traditional OCR-based methods, further reinforcing the advantages of vision-language pretraining.
5.0.0.0.3 Language Tokenization
Tokenization methods, such as Byte Pair Encoding (BPE) Shibata et al. (1999); Sennrich et al. (2016), are widely used in language modeling, but recent studies suggest that they may not always be optimal. For instance, MegaByte Yu et al. (2023) demonstrated that fixed-length tokenization can improve both computational efficiency and cross-modal capabilities. Similarly, BLT Pagnoni et al. (2024) proposed entropy-based tokenization, while LCM team et al. (2024) emphasized the benefits of processing higher-level semantic concepts rather than individual tokens. Inspired by these approaches, we explore whether adaptive image patches can effectively infer textual meaning. At a higher level, we investigate the unification of text and image inputs into a shared representation space, enabling reasoning through abstract semantic concepts rather than traditional token-based methods.
@@ 6 Conclusion
We present PixelWorld, a benchmark that renders text, tables, code, and images as pixels, enabling direct evaluation of the Perceive Everything as Pixels (PEAP) paradigm. Experiments yield three takeaways: (1) Semantic understanding. PEAP matches token baselines on sentence-/paragraph-level tasks, while its patch-level attention closely mirrors token attention, pointing toward “vision-as-token” models. (2) Reasoning. Accuracy drops on math, logic, and program-repair benchmarks; Chain-of-Thought prompts narrow but do not eliminate this gap. (3) Multimodal tasks. Pixel input outperforms OCR pipelines on websites, slides, and documents by preserving spatial context and avoiding recognition errors. To curb the higher latency of pixel inputs, we introduce PEAP-Fast, which prunes blank patches and accelerates inference by up to $3\times$3 × without hurting accuracy. Taken together, these findings underscore both the promise and trade-offs of PEAP. PixelWorld thus serves as a practical stress test and diagnostic benchmark, encouraging the community to adopt PEAP as a holistic yardstick while guiding research on efficiency improvements and on closing the reasoning gap in next-generation multimodal agents.
Limitations
Due to constraints in time and computational resources, we could not comprehensively evaluate all benchmarks and language models. To simplify our experimental setup, we primarily focused on tasks that can be evaluated through straightforward string-based computations. Tasks involving complex settings such as function calling, code reasoning, or retrieval remain insufficiently covered.
Furthermore, unless prompts were explicitly provided by the dataset, we employed uniform Direct Prompting and Chain-of-Thought prompting strategies, which may not be representative enough and could introduce biases. More sophisticated prompting techniques, such as ReAct, were not thoroughly evaluated.
Finally, our framework focuses specifically on evaluating the textual reasoning abilities of vision-language models, leaving other multimodal models—such as video generation models that lack textual reasoning capability—beyond the scope of this evaluation.
Ethics Statement
This paper focuses on the reasoning capabilities and interpretability of vision-language models. Vision-language models possess strong abilities for understanding and interacting with the real world, and thus carry potential risks of misuse. However, our methods concentrate on benign reasoning tasks and propose interpretability mechanisms that can help mitigate potential misuse.
Additionally, the datasets used in our experiments primarily comprise English data. While our framework is potentially transferable to other languages, vision-language models may suffer from reduced performance due to limited availability of resources in languages other than English.
Reproducibility Statement
All code used in this paper has been made publicly available, along with intermediate results of all synthetic data. Additionally, we set the inference temperature uniformly to 0 across all models. Thus, we believe our work is highly reproducible. However, minor discrepancies may arise due to differences in hardware and inference frameworks, potentially leading to results that are not strictly identical.
@@ Appendix B Attention Heatmap before and after ImageFast Method
Appendix C Model Size and Computational Budget
All data syntactic process and experiments used around 50 hours of a 32-core CPU host and around 200 GPU hours on Nvidia A6000. The total cost of API calls was less than 1,000 USD.
Appendix D Usage of AI Assistants
During the development process, we drafted the code framework and data structures by ourselves initially, using Copilot to assist in enhancing development efficiency. Additionally, in the writing phase, we only employed ChatGPT for proofreading purposes, such as correcting grammatical errors and properly formatting tables.